what is the axis of symmetry for the quadratic equation: h(x)= 1-1/2t-t^2
Answer by lwsshak3(11628)

what is the axis of symmetry for the quadratic equation: h(x)= 1-1/2t-t^2
h(t)=-t^2-(1/2)t+1
complete the square:
h(t)=-(t^2+(1/2)t+1/16)+1/16+1
h(t)=-(t+(1/4))^2+17/16
This is a parabola that opens downward with vertex at (-1/4,17/16)
axis of symmetry:x=-1/4